OpenStudy (anonymous):

find slope of the line passing through the points (2,5) and (8,-4)

OpenStudy (anonymous):

please show with steps

OpenStudy (shaik0124):

slope =y2-y1/x2-x1 let us take x1,y1 =2,5 x2,y2=8,-4 now substitute this values

OpenStudy (anonymous):

ok